OpenStudy (anonymous):

It does not matter which of the two points of a line you choose to call (x1, y1) and which you choose to call (x2, y2) to calculate the slope of the line.

OpenStudy (texaschic101):

true... example : (5,3) x1 = 5 and y1 = 3 (3,2) x2 = 3 and y2 = 2 slope(m) = (y2 - y1) / (x2 - x1) slope(m) = (2 - 3) / (3 - 5) slope(m) = -1/-2 = 1/2 ==> slope is 1/2 now lets switch our points (3,2) x1 = 3 and y1 = 2 (5,3) x2 = 5 and y2 = 3 slope(m) = (y2 - y1) / (x2 - x1) slope(m) = (3 - 2) / (5 - 3) slope(m) = 1/2 ==> slope is 1/2

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

your welcome :)